OpenStudy (anonymous):

Sarra had 7/8 pounds of turkey. She used 2/3 pounds of the turkey for sandwiches. How many pounds of turkey does she have left? A. 5/24 pounds B. 1/4 pounds C. 5/12 pounds D. 5/8 pounds @prettyboy937 @PaulaLovesSchool13 @baking4lyf @iPurple @monsterfox1997 please help fan and medal!!!

OpenStudy (anonymous):

How many pounds of turkey does she have left? hat operation will we do?

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

what*

OpenStudy (anonymous):

@prettyboy937 you can't do that! http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

you gave a direct answer

OpenStudy (anonymous):

no oh well, that is cheating and wrong

OpenStudy (igreen):

^^

OpenStudy (anonymous):

Meg needs a helpful answer to help next time

OpenStudy (anonymous):

@Megzz2002 tell them its ok

OpenStudy (igreen):

\(\dfrac{7}{8} - \dfrac{2}{3} =~?\) @Megzz2002

OpenStudy (anonymous):

Meg, you know its wrong

OpenStudy (igreen):

@prettyboy937 Do you want me to tag a Moderator and ask if it's okay?

OpenStudy (anonymous):

^

OpenStudy (anonymous):

be my guess

OpenStudy (anonymous):

guess?

OpenStudy (anonymous):

Green, tag them

OpenStudy (anonymous):

I don't know I really just want to understand what I'm supposed to be doing

OpenStudy (igreen):

@ganeshie8 @Callisto @shrutipande9

OpenStudy (anonymous):

See, Meg, needs HELP

OpenStudy (igreen):

Yes, and you can't do that if @prettyboy937 is giving answers.

OpenStudy (anonymous):

well then don't tag me in any more questions bye

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

I am finding a link for you Meg

OpenStudy (anonymous):

my parents said to go on openstudy so I could get help with my questions and learn how to do the equations myself

OpenStudy (anonymous):

And you won't learn if someone just pops a anser at you

OpenStudy (anonymous):

Here is a link to subtract fractions http://www.mathsisfun.com/fractions_subtraction.html Math is fun is a website with soooo many math topics for you

OpenStudy (igreen):

To find the answer to this problem, you have to subtract the two fractions: \(\dfrac{7}{8} - \dfrac{2}{3}\) First, you have to find a common denominator, which is 24 here. Now convert the fractions into denominators with 24: \(\dfrac{7 \times 3}{8 \times 3} - \dfrac{2 \times 8}{3 \times 8} \rightarrow \dfrac{21}{24} - \dfrac{16}{24}\) Now can you subtract 21 - 16? @Megzz2002

OpenStudy (anonymous):

thanks @PaulaLovesSchool13

OpenStudy (anonymous):

Okay now listen to Mr. Green

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Uh.....green?

OpenStudy (anonymous):

@Megzz2002 u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

am I right?

OpenStudy (igreen):

Yes, that gives us \(\dfrac{21}{24} - \dfrac{16}{24} \rightarrow\dfrac{5}{24}\).

OpenStudy (igreen):

See? @prettyboy937 answer was wrong too. Nothing worse than a WRONG DIRECT ANSWER.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks soo much @iGreen and @PaulaLovesSchool13

OpenStudy (anonymous):

Meg, make sure you check out the link I sent ;)

OpenStudy (anonymous):

oh sure :D

OpenStudy (anonymous):

why tf are you still talking to me

OpenStudy (anonymous):

what the elephants, you are still in here? o_o

OpenStudy (anonymous):

I did check out the link you posted thanks !!!:)

OpenStudy (igreen):

I just tagged him. ;) @PaulaLovesSchool13

OpenStudy (anonymous):

oh thanks a lot Green -_- he is angryyyyy And that is great Meg!